In coming to a stop, a motorcycle leaves skid marks 100m long on the highway. Assume that the motorcycle slows down at a constant rate of -7.5m/s2, estimate the speed of the car just before breaking.


1
Expert's answer
2020-10-21T06:28:39-0400

We can find the speed of the motorcycle just before breaking from the kinematic equation:


"v^2=v_0^2+2as,"

here, "v=0\ \dfrac{m}{s}" is the final velocity of the motorcycle, "v_0" is the initial velocity of the motorcycle (just before breaking), "a=-7.5\ \dfrac{m}{s^2}" is the deceleration of the motorcycle and "s=100 \ m" is the distance traveled by the motorcycle when it is breaking.

From this formula we can calculate the speed of the motorcycle just before breaking:


"v_0=\sqrt{-2as}=\sqrt{-2\cdot(-7.5\ \dfrac{m}{s^2})\cdot 100\ m}=38.73\ \dfrac{m}{s}."

Answer:

"v_0=38.73\ \dfrac{m}{s}."